Thermal Control Procedures

  Revised 13 July 2020 by A. R. Walker.  
Original by Maxime Boccas, 18 October 1999.

Over the period 1993-2000, many improvements have been carried out at the Víctor M. Blanco 4-meter Telescope to improve its image quality. The principal changes are: an active primary mirror support with lookup tables, a refigured f/8 secondary mirror, an image analyzer for the Cassegrain foci, removal of most heat sources inside the dome, control of the dome temoerature during the day using the AHUs and heicopter fan, active ventilation (air sucker) of the primary cell at night, dome ventilation doors and active control of oil temperature.  The dome outer surface is covered with insulated aluminum panels which although the outer surface heats up during the day mpre than the originbal Lomit paint, does not over-cool at night..

Since 2012, DECam at Prime Focus has introduced a 5-axis hexapod (x,y,z, tip, tilt) position of the instrument and corrector, driven by a Look Up Table and tweaked on an exposure by exposure basis by a wavefront sensor that analyses out-of-focus inages, and makes a "tweak" to the hexapod position.

1. Weather and thermal issues

The night assistant must check that all the environmental control functions explained below are working properly during the night.

1.1 General closure conditions:

ABSOLUTE closure when:

  • wind > 45 mph on average over the last 30 minutes (the telescope might still not shake badly but the air is dusty and will bring too much dirt inside the domes)
  • If wind > 25 mph absolute closure of the lateral sliding doors
  • humidity > 85 %, as indicated by the met tower 
  • Tmirror - Tdew < 2°C. The values of Tdew and Tmirror-Tdew are shown in Blanco ECS:Sensors Diagram, which is actualized every 5 minutes.
  • See calculating risk of condensation for more information about dew point.

Dome re-opening:

  • if wind <= 45mph on average over the last 30 minutes
  • if RH < 85% for 30 consecutive minutes (in all cases, check outside that no water is dripping from the buildings).
  • if Tmirror - Tdew >= 2°C for 30 consecutive minutes


  1.  If  wind > 25 mph, avoid observing into the wind direction as it could jitter the telescope and will bring dirt inside the dome.
  2.  If wind > 15 mph   evaluate image quality ,close doors if need be, raise wind blind ( A.K.A.: windscreen ) 

BEWARE PLEASE: Don't permit observing beyond these limits! If you have trouble to be heard by the astronomer, please REPORT it (so we can politely explain the rules).

Back to Top

1.2. Dome opening:

The dome AND the mirror cover have to be opened 1 hour before sunset, orienting the shutter eastward to make sure no direct sunlight will enter the dome. Dome and mirror cover have to be closed at the end of the night.

1.3. Lateral sliding doors opening:

These doors have to be opened at sunset and left opened all night.

  • If the wind > 15 mph on average, evaluate the image quality to decide whether to close them or not (elongated images mean you have to close the doors).
  • If wind > 25 mph on average, ABSOLUTE closure (because of wind-shake on the telescope and deposit of dust).

1.4 Primary mirror cooling 

The Primary mirror cooling system has from 2019 been activated only in "sucking" mode, all the time, day and night. That is,we do no longer blow cold air onto the primary.        

  • Let's recall that a 1° temperature difference between the mirror surface and the ambient air will create additional mirror-seeing (typically up to 0.5" extra for 1° difference). Thus our goal is to maintain the mirror temperature within 1° of the predicted temperature of the following night.

Back to Top

During the night, you are invited to check  Blanco ECS Telemetry and check that:

  • the 4 sensors of the mirror are within a 0.5-1° range
  • the mirror temperature, dome temperature and outside temperature are within a 1° range
  • Any problems detected should be reported by email

1.5. Floor cooling, Dome Air Handling Units (AHUs), and stirring fan on main floor (M):

The cooling of the C floor and M floor in the dome is always ON. It is currently regulated automatically.

The two AHUs should always be off when the dome is open.  There is no automatic control,  Thus the procedure is for the telescope operator to turn green on at the end of the night, and then Telops staff evaluate the inside versus outside temperature and the weather forecast prediction for the following night, then decide whether to turn off one or both, and at what time.

BEWARE:  In conditions of high humidity the AHUs can ice up and restrict the air flow,   This condition can be cleared rapidly by turning off the cooling, air flow on.

The large stirring fan (helicopter) on the M floor (inside the dome) should always be ON when the dome is closed and OFF when the dome is opened. This fan improves greatly the air circulation inside the dome.

SAFETY:   Both the AHUs and the helicopter can be turned off during instrument changes or any other works around the telescope if the noise is a safety concern, e.g. by impeding communication.  But don't forget to turn them back on when you are done!

Back to Top

1.6. Oil cooler:

The cooling of the oil circulating under the R.A. rear and front pads is activated automatically when the pumps are turned ON by Observer support at sunset. The oil glycol valve is controlled (range is 1-2.25 V) by an equation involving the low dome and oil glycol temperature sensors. It is important that the oil cooling be working: if it is not, you can diagnostic it in ECS Telemetry (the "before pad" and "glycol" entries on the OIL line will quickly indicate around 30°).

The oil temperature for the Hidrostatic Bearings must be between 8 to 10C.


1.7. Chimney fans:   Not in use.

These fans are located inside the chimney at the level of the primary mirror and suck ambient air into the mirror cell so as to form a laminar downward flow in the chimney. This is to prevent Cass cage heat sources from creating convection in the light path in the chimney.

At Cassegrain foci (f/8), Observer Support has to turn ON the chimney fans before sunset. The switch is on the power supply on the old console room desk. The night assistant must turn it OFF at the end of the night.

At Prime Focus:   The top of the chimnney is covered with a cap.

1.8. Air conditioning:

The current thermal plan in the building is not changed. Always leave the air conditioning at full power. Do not use heat sources at level M, MZ and C. Always maintain the doors closed in the passageways (especially the glass door in front of the lift at GR floor, the cryocooler compressor and pump rooms). In general, respect the signs in the building.

Back to top


2. Brief reminder of known optical problems

The image quality of the telescope can suffer deterioration mainly from 3 optical problems:

Note IMAN is normally used only on engineering nights.

2.1. Astigmatism

 This is due to tiny slippage of M1 and/or deformation under its own gravity. Astigmatism will show up as elongated images, perpendicular on each side of the focus, especially for large H.A. (typically more than 2 hours). We compensate that with a lookup table, which controls the pressure in the air bags under M1. Thus the TCP toggle "CORR" should always be ON, at all foci. Note that if the F/8 secondary mirror lose its vacuum, strong astigmatism will immediately show up (typically 2-4 microns as measured by iman).

2.2. Coma

This aberration shows up in 2 cases: at Prime Focus if the primary mirror axis is not aligned with the instrument and in the 2-mirror configuration (f/8) when there is a misalignment between them. Coma produces images with a flare (a tail like a comet) which is identical (same amplitude and direction) on each side of the focus.

At Prime Focus coma is controlled by the DECan active optics system and Look Up Table that control the hexapod on which the instrument is mounted.

Coma usually increases the further you go from the optical axis -this is called "field coma" and is normal (it is always supposed to be 0 on axis)- and also shows up when the optical axis of M1 and M2 are not coaligned -this is called "decentering coma" and is abnormal-. Our f/8 is a Ritchey-Chretien (RC) type, which means it is optically designed to correct the field coma. If the telescope is properly aligned there should not be any visible coma at f/8. If coma shows up anywhere in the field (it would usually be constant all over the field), there is a mirror misalignment and this can be corrected by using IMAN.

Back to Top

2.3. Defocus

Thermal expansion that can occur at night will cause the telescope to defocus. The prime focus changes at a rate of -110 microns/°C, the f/8 focus at -780microns/°C. The actual movement of He f/8 mirror is a factor 10 less (i.e. -78 microns/°C).  The auto[focus of DECam means you do not have to worry about focus at prime focus.

Use Blanco ECS Sensors Diagram to check the Serrurier truss temperature and adjust the focus accordingly (F/8).

3. Prime Focus nights:

 Focus decreases by 110 units per 1° increase in temperature.

Back to Top

4. f/8 nights: 

* CORR ON all the time.

* M2 Corrections ON all the time 

* Temperature drift compensation:

Focus number increases when focal plane (and M2) moves up

The focus number represents microns of motion of the focal plane (not of M2)

Decrease focus setting 780 units per 1° increase in temperature


*Collimation:this has to be checked according to the following procedure:

  • On the first night of a f/8 run, check in the TCS the tilt value (amplitude and azimuth) written on the white board and in the logbook. If this is not the case, choose /Absolute_Tilt and enter the right values. Next, run IMAN according to the instructions in the IMAN instruction book (a version is also available on the web). Write the new tilt values on the white board and in the logbook.

Back to Top